OpenStudy (anonymous):

solve (6/t)-2/(t-1)=1

OpenStudy (anonymous):

so im guessing you have to solve for t?

OpenStudy (anonymous):

welcome to OS btw

OpenStudy (anonymous):

\[(6-t) - \frac{ 2 }{ (t-1) } = 1\] \[\frac{ (6-t)*(t-1) - 2 }{ (t-1) } = 1\]

OpenStudy (anonymous):

\[\frac{ (6t-t^2-6 - t) - 2 }{ (t-1) } = 1\] \[\frac{ (5t-t^2-8) }{ (t-1) } = 1\] \[ (5t-t^2-8) = 1 *(t-1)\] \[ (5t-t^2-8) = t-1\] \[ 5t-t^2-8 - (t-1) = 0\] \[ 4t-t^2-7 = 0\] \[ t^2- 4t + 7 = 0\]

OpenStudy (anonymous):

solve the quadratic